the second of 2 numbers is 7 times the first. their sum is 32. find the numbers
Answer by oberobic(2304)

Two numbers: x and y.
y = 7x
.
x + y = 32
so
y = 32-x
.
substitute
.
7x = 32-x
8x = 82
x = 4
.
y = 32-4 = 28
.
Is the second number 7 * first number?
28 = 7(4)
Correct.
.
Answer: The numbers are 4 and 28.
.
Done.